# WildBench: Benchmarking LLMs with Challenging Tasks from Real Users in the Wild

### v1

@@ Abstract
We introduce WildBench, an automated evaluation framework designed to benchmark large language models (LLMs) using challenging, real-world user queries. WildBench consists of 1,024 tasks carefully selected from over one million human-chatbot conversation logs. For automated evaluation with WildBench, we have developed two metrics, WB-Reward and WB-Score, which are computable using advanced LLMs such as GPT-4-turbo. WildBench evaluation uses task-specific checklists to evaluate model outputs systematically and provides structured explanations that justify the scores and comparisons, resulting in more reliable and interpretable automatic judgments. WB-Reward employs fine-grained pairwise comparisons between model responses, generating five potential outcomes: much better, slightly better, slightly worse, much worse, or a tie. Unlike previous evaluations that employed a single baseline model, we selected three baseline models at varying performance levels to ensure a comprehensive pairwise evaluation. Additionally, we propose a simple method to mitigate length bias, by converting outcomes of “slightly better/worse” to “tie” if the winner response exceeds the loser one by more than $K$ characters. WB-Score evaluates the quality of model outputs individually, making it a fast and cost-efficient evaluation metric. WildBench results demonstrate a strong correlation with the human-voted Elo ratings from Chatbot Arena on hard tasks. Specifically, WB-Reward achieves a Pearson correlation of 0.98 with top-ranking models. Additionally, WB-Score reaches 0.95, surpassing both ArenaHard’s 0.91 and AlpacaEval2.0’s 0.89 for length-controlled win rates, as well as the 0.87 for regular win rates.
@@ 1 Introduction
Large language models (LLMs) have become integral to a wide range of real-world applications due to their strong generalization capabilities across diverse tasks. However, effectively evaluating their performance remains a challenging problem, particularly when striving for an automated and cost-effective solution. Traditional benchmarking datasets like MMLU [13] focus primarily on assessing the reasoning abilities of LLMs using multiple-choice questions, which fall short in evaluating the more open-ended problems that real-world users pose. Chatbot Arena [4] provides an online platform where human preferences are collected to judge pairs of model outputs, subsequently ranking LLMs using Elo ratings. While this human-based evaluation method offers valuable insights into user preferences, it has notable limitations, such as high labor costs, the inability to deliver real-time results, a lack of data transparency, and the challenge of fairly evaluating all models with same data.
@@
Several automated benchmarks such as AlpacaEval [15], MT-bench [38], and ArenaHard [14] employ advanced LLMs like GPT-4-Turbo to assess the quality of model responses. Comparative analyses of these benchmarks are presented in Table 1 and Figure 3. These existing benchmarks exhibit significant shortcomings in task composition and skill coverage, particularly in mirroring the natural distribution of real-world user tasks. MT-bench, comprising only 80 hand-crafted examples, lacks sufficient breadth for a comprehensive evaluation. Meanwhile, AlpacaEval, with 805 tasks derived from multiple alignment datasets, includes relatively simple tasks, such as “What is the capital of Australia?” and suffers from low task diversity; for instance, over 20 tasks redundantly assess recipe generation skills. This benchmark mostly focuses on information-seeking tasks, containing merely 6% coding and 3% mathematics tasks. Conversely, ArenaHard, sampling 500 tasks from ChatbotArena, displays an excessive concentration on coding and debugging tasks, accounting for over 57% of its content. Most existing benchmarks do not sufficiently challenge the models with varied and unexpected nature of user inquiries in practical settings, thus limiting their overall effectiveness in providing a holistic evaluation. This issue highlights the necessity for more comprehensive benchmarks that can better simulate the wide range of tasks from real users.
@@
In this paper, we introduce WildBench, an automated evaluation framework engineered to assess LLMs using complex queries from real-world users. The examples in WildBench are periodically updated, with the current version (V2) comprising 1,024 tasks carefully curated from real user-chatbot dialogs provided by the WildChat project [37]. We engage multiple advanced LLMs to process a filtered selection from WildChat, tasking them with the analysis of the requisite knowledge and skills for each task and subsequently labeling the difficulty level. Tasks considered as easy by all models are excluded. We ensure the distribution of tasks mirrors the original WildChat data, such that the task distribution of WildBench is still natural (Figure 3). Additionally, all finalized tasks undergo manual review. Further details are provided in Section 2.
Figure 1: Evaluation framework for WildBench.
@@
WildBench evaluation is illustrated in Figure 1. To design a reliable automatic evaluation, we employ two key designs for prompting LLMs as judges. Drawing inspiration from how humans evaluate responses to open-ended questions, we develop task-specific checklists. These checklists guide LLMs in generating consistent and reliable judgments, with each checklist comprising questions focused on specific criteria. Similar to the zero-shot Chain-of-Thoughts (CoT) prompting [12], we prompt LLMs to provide step-by-step, structured analyses of each LLM response. This method encourages a detailed, fine-grained evaluation process, culminating in a well-justified final decision.
@@
To report performance, we employ two primary metrics: WB-Reward for pairwise comparisons and WB-Score for individual scoring. WB-Reward is based on pairwise comparisons between LLMs, with five possible outcomes: “A is much/slightly better/worse than B” or “Tie.” Notably, we used three baseline models to compare with each testing model instead of using a single baseline model, as most prior works do. This approach provides a more comprehensive assessment based on different levels of model performance. WB-Score measures the quality of each model’s generation individually, offering a quicker and more cost-effective evaluation. To mitigate the bias towards longer outputs, a common issue in LLM-as-a-judge evaluations [6], we introduced a simple length-penalty method, converting slight wins/losses to ties when the winner’s output is significantly longer than the loser’s.
@@
Both metrics have demonstrated strong correlations with human judgments, evidenced by a Pearson correlation of 0.98 for WB-Reward and 0.95 for WB-Score against the human-voted Elo rating from Chatbot Arena on the top-ranking models. These scores significantly surpass other benchmarks, such as ArenaHard[14]’s 0.91 and AlpacaEval2.0’s 0.87 (0.89 for LC)[15, 6], validating WildBench’s effectiveness and alignment with human-based evaluation. More details are shown in Tab 3 in Sec 4.
@@ 2 WildBench Data Curation
Table 1: Statistical comparison of LLM alignment benchmarks.
@@ 2.1 Mining Challenging Tasks from WildChat
We sourced tasks from the WildChat dataset [37], which comprises one million human-chatbot conversations from real users. [^1] WildChat is released under the AI2 ImpACT license. This dataset is particularly suited for conversion into an evaluation benchmark because it contains a diverse array of tasks that users expect LLMs to perform, such as writing assistance, coding, mathematics, data analysis, role playing, and planning.
@@
To control the quality and diversity of the selected tasks, we applied several filtering steps. First, we removed user queries that were either too short (less than 10 tokens) or excessively long (more than 3,000 tokens). We also excluded conversations with more than five user-chatbot turns to maintain focus and coherence in the tasks, as conversations exceeding five turns tend to contain multiple topics. Furthermore, we focused on English data and filtered out non-English tasks. Since our focus is more on evaluating the capabilities of LLMs rather than content moderation, we also removed toxic conversations. To ensure task diversity, we used sentence embeddings from SentenceBERT [30] to calculate the cosine similarity between queries, discarding those with a high similarity score above 0.9. The threshold is determined by manual inspection. Lastly, to further enhance task diversity, we used a diverse user pool by retaining only the last conversation for each unique device, thus removing tasks from the same user that might require similar underlying skills.
@@
To identify challenging tasks that can distinguish the performance of different LLMs, we used GPT-4-Turbo [26], Claude-3-Sonnet, and Opus [2] to analyze the required background knowledge and reasoning capabilities for each task. These models assigned a difficulty rating on a five-point scale (from “very easy,” “easy,” “medium,” “hard,” to “very hard”). Tasks rated as “very easy” or “easy” by all models were excluded. From the remaining pool, we randomly sampled 1,500 tasks to ensure that the distribution of task categories is similar to the original dataset.
@@
To improve the quality of selected tasks, human annotation was used for quality control. We first used GPT-4-Turbo to summarize the intent of each query. These summaries were then reviewed to remove nonsensical tasks. Finally, we retained 1,024 tasks for WildBench .
@@ 2.2 WildBench Statistics
Table 1 compares the statistics of WildBench to existing benchmarks AlpacaEval [15, 6], MT-Bench [38], and ArenaHard [14]. Among these benchmarks, only ArenaHard and WildBench are sourced from user queries in the wild (“RealUser”), rather than being curated by experts or through crowdsourcing. The difference between ArenaHard and our WildBench is that our data distribution aligns with real users’ task categories, rather than overly focusing on coding and debugging as ArenaHard does.
@@
To enable a fine-grained analysis of LLM capabilities across varied tasks, we categorize the tasks into 12 categories based on previous analysis of ShareGPT queries [27] and our intent annotation of the tasks. [^2] Detailed descriptions about the 12 task categories are shown in Appendix E. The distribution of the task categories is shown in Figure 3. In this figure, we also compare to AlpacaEval and ArenaHard. Notably, WildBench is more balanced compared to AlpacaEval and ArenaHard, which have over 50% of their tasks in Information seeking and Coding & Debugging categories, respectively.
@@ 3.1 Instance-Specific Checklists
Powerful LLMs have been widely used as judges to evaluate the quality of LLM outputs in many automatic evaluation methods, such as AlpacaEval [15]. However, even asking humans to judge which of the given two model outputs is better can be subjective and ambiguous. Moreover, such judgements provide limited information about the quality of the models. Without a constant, interpretable, and comprehensive evaluation standard, the results can be noisy and hard to interpret.
To address this issue, we generate a checklist for each test query in WildBench to comprehensively evaluate the responses of different models. The checklist consists of 5-10 questions that are designed to be interpretable and easy to verify. We combine the responses of GPT-4-Turbo and Claude-3-Opus to finalize the checklists, thereby mitigating the bias of using a single LLM as the evaluator. These checklists have been manually reviewed and are used as part of the prompts for LLM judges to evaluate the responses of different models. An example of the checklist can be found in Figure 1.
@@ 3.2 Pairwise Evaluation with WB-Reward Metric
In Figure 1, we detail the step-by-step evaluation process for pairwise comparison. First, we provide a chain of evaluation questions to guide the LLM judge to analyze the user query and the conversation history. The LLM then evaluates the two responses and also analyze where and why one is better than the other. Finally, we ask the LLM to make a final judgment on which response is better and why. This method is inspired by the evaluation process in human evaluation, where human judges are asked to provide detailed feedback on the quality of the responses before making a final decision. The full evaluation prompt can be found at Appendix A
@@
Using a single baseline model for pairwise evaluation can lead to noisy and biased evaluations. To mitigate this issue, we use three baseline models (GPT-4-Turbo-0429, Claude-3-Haiku, and Llama-2-70B-chat [34]) to compute the rewards for each model. Our metric WB-Reward (Mix) is the average of the rewards from these three baselines on 1024 examples, providing a more robust performance evaluation on WildBench.
@@
Previous studies have shown that LLM judges tend to prefer longer responses [6]. To mitigate this bias, we propose a simple and intuitive length penalty method. If the winning response is longer than the losing one by a certain threshold (K characters), we convert Slightly Win/Slightly Lose to a Tie. $K$ can be customized via our leaderboard web-page for personalized configuration. Setting $K=\infty$ will disable the length penalty.
@@ 3.3 Individual Evaluation with WB-Score Metric
Although pairwise evaluation provides a direct comparison between LLMs, it is usually more expensive and time-consuming than grading each individual LLM generation. To individually evaluate the performance of each model on WildBench, we prompt GPT-4-Turbo to assign a score from 1 to 10 for each model’s response. The full evaluation prompt can be found at Appendix B.
@@
4 Results and Analysis
In this section, we analyze the performance of different models on WildBench. We first present the leaderboard analysis, then examine the length bias issue in the evaluation process, and finally discuss the correlation between WildBench-Score and ChatbotArena Elo rating.
Table 2: Evaluation results of LLMs using WildBench and other benchmarks. Please refer to Figure 5 and our website linked here to view and interact with the full results.

### v2

@@ Abstract
We introduce WildBench, an automated evaluation framework designed to benchmark large language models (LLMs) using challenging, real-world user queries. WildBench consists of 1,024 examples carefully selected from over one million human-chatbot conversation logs. For automated evaluation with WildBench, we have developed two metrics, WB-Reward and WB-Score, which are computable using advanced LLMs such as GPT-4-turbo. WildBench evaluation uses task-specific checklists to evaluate model outputs systematically and provides structured explanations that justify the scores and comparisons, resulting in more reliable and interpretable automatic judgments. WB-Reward employs fine-grained pairwise comparisons between model responses, generating five potential outcomes: much better, slightly better, slightly worse, much worse, or a tie. Unlike previous evaluations that employed a single baseline model, we selected three baseline models at varying performance levels to ensure a comprehensive pairwise evaluation. Additionally, we propose a simple method to mitigate length bias by converting outcomes of “slightly better/worse” to “tie” if the winner’s response exceeds the loser’s by more than $K$ characters. WB-Score evaluates the quality of model outputs individually, making it a fast and cost-efficient evaluation metric. WildBench results demonstrate a strong correlation with the human-voted Elo ratings from Chatbot Arena on hard tasks. Specifically, WB-Reward achieves a Pearson correlation of 0.98 with top-ranking models. Additionally, WB-Score reaches 0.95, surpassing both ArenaHard’s 0.91 and AlpacaEval2.0’s 0.89 for length-controlled win rates, as well as the 0.87 for regular win rates.
@@ 1 Introduction
Large language models (LLMs) have become integral to a wide range of real-world applications due to their strong generalization capabilities across diverse tasks. However, effectively evaluating their performance remains a challenging problem, particularly when striving for an automated and cost-effective solution. Traditional benchmarking datasets like MMLU (Li et al., 2023a) focus primarily on assessing the reasoning abilities of LLMs using multiple-choice questions, which fall short in evaluating the more open-ended problems that real-world users pose. Chatbot Arena (Chiang et al., 2024) provides an online platform where human preferences are collected to judge pairs of model outputs, subsequently ranking LLMs using Elo ratings. While this human-based evaluation method offers valuable insights into user preferences, it has notable limitations, such as high labor costs, the inability to deliver real-time results, a lack of data transparency, and the challenge of fairly evaluating all models with the same data.
@@
Several automated benchmarks such as AlpacaEval (Li et al., 2023b), MT-bench (Zheng et al., 2024), and ArenaHard (Li et al., 2024) employ advanced LLMs like GPT-4-Turbo to assess the quality of model responses. Comparative analyses of these benchmarks are presented in Table 1 and Figure 3. These existing benchmarks exhibit significant shortcomings in task composition and skill coverage, particularly in mirroring the natural distribution of real-world user tasks. MT-bench, comprising only 80 hand-crafted examples, lacks sufficient breadth for a comprehensive evaluation. Meanwhile, AlpacaEval, with 805 tasks derived from multiple alignment datasets, includes relatively simple tasks, such as “What is the capital of Australia?” and suffers from low task diversity; for instance, over 20 tasks redundantly assess recipe generation skills (e.g., “can you provide a recipe for …?”). We show a few examples in Figure 1 to illustrate the differences between AlpacaEval and our WildBench.
AlpacaEval mostly focuses on information-seeking tasks, containing merely 6% coding and 3% mathematics tasks. Conversely, ArenaHard, sampling 500 tasks from ChatbotArena, displays an excessive concentration on coding and debugging tasks, accounting for over 57% of its content. Most existing benchmarks do not sufficiently challenge the models with the varied and unexpected nature of user inquiries in practical settings, thus limiting their overall effectiveness in providing a holistic evaluation. This issue highlights the necessity for more comprehensive benchmarks that can better simulate the wide range of tasks from real users.
@@
In this paper, we introduce WildBench, an automated evaluation framework designed for assessing LLMs using complex tasks from real-world users. The examples in WildBench are periodically updated, with the current version (V2) comprising 1,024 tasks carefully curated from real user-chatbot dialogs provided by the AI2’s WildChat project (Zhao et al., 2024). We engage multiple advanced LLMs to process a filtered selection from WildChat, tasking them with the analysis of the requisite knowledge and skills for each task and subsequently labeling the difficulty level. Tasks considered as easy by all models are excluded. We ensure the distribution of tasks mirrors the original WildChat data, such that the task distribution of WildBench is still natural (Figure 3). Additionally, all finalized tasks undergo manual review. Further details are provided in Section 2.
As shown in Figure 1, WildBench presents a significantly harder challenge due to the complexity, depth, and realism of the tasks involved. WildBench is sourced from real-world user interactions and has been carefully curated to ensure diversity and challenge. The tasks in WildBench typically demand higher-order reasoning, such as writing and/or debugging code with specific constraints, creative writing with multiple constraints on the style and content, or designing a software system with complex requirements. These tasks often require critical thinking, creativity, and technical expertise, making WildBench substantially more challenging than AlpacaEval, where simpler, factual, or surface-level tasks dominate.
Figure 1: Example tasks sampled from AlpacaEval (Li et al., 2023b) and WildBench. Tasks in WildBench are more diverse and challenging, which are collected from real users in the wild. Complex real-user tasks usually have multiple constraints and require higher-order reasoning skills, which are well represented in WildBench.
@@
WildBench evaluation is illustrated in Figure 4. To design a reliable automatic evaluation, we employ two key designs for using LLMs as judges. Drawing inspiration from how humans evaluate responses to open-ended questions, we develop task-specific checklists. These checklists guide LLMs in generating consistent and reliable judgments, with each checklist comprising questions focused on specific criteria. Similar to the zero-shot Chain-of-Thoughts (CoT) prompting (Kojima et al., 2022), we prompt LLMs to provide step-by-step, structured analyses of each LLM response. This method encourages a detailed, fine-grained evaluation process, culminating in a well-justified final decision.
@@
We employ two primary metrics: WB-Reward for pairwise comparisons and WB-Score for individual scoring. WB-Reward is based on pairwise comparisons between LLMs, with five possible outcomes: “A is much/slightly better/worse than B” or “Tie.” Notably, we used three baseline models to compare with each testing model instead of using a single baseline model, as most prior works do. This approach provides a more comprehensive assessment based on different levels of model performance. WB-Score measures the quality of each model’s generation individually, offering a quicker and more cost-effective evaluation. To mitigate the bias towards longer outputs, a common issue in LLM-as-a-judge evaluations (Dubois et al., 2024), we introduced a simple length-penalty method, converting slight wins/losses to ties when the winner’s output is significantly longer than the loser’s.
@@
Both metrics have demonstrated strong correlations with human judgments, evidenced by a Pearson correlation of 0.98 for WB-Reward and 0.95 for WB-Score against the human-voted Elo rating from Chatbot Arena on the top-ranking models. These scores significantly surpass other benchmarks, such as ArenaHard(Li et al., 2024)’s 0.91 and AlpacaEval2.0’s 0.87 (0.89 for the length-controlled version) (Li et al., 2023b; Dubois et al., 2024), validating WildBench’s effectiveness and alignment with human-based evaluation. More details are shown in Table 3 in Section 4.
@@ 2 WildBench Data Curation
Table 1: Statistical comparison of LLM alignment benchmarks. Length are in characters.
@@ 2.1 Mining Challenging Tasks from WildChat
We sourced tasks from the WildChat dataset (Zhao et al., 2024), which comprises one million human-chatbot conversations from real users. This dataset is particularly suited for conversion into an evaluation benchmark because it contains a diverse array of tasks that users expect LLMs to perform, such as writing assistance, coding, mathematics, data analysis, role playing, and planning.
@@
To control the quality and diversity of the selected tasks, we applied several filtering steps. First, we removed user queries that were either too short (less than 10 tokens) or excessively long (more than 3,000 tokens). We also excluded conversations with more than five user-chatbot turns to maintain focus and coherence in the tasks, as conversations exceeding five turns tend to contain multiple topics. Furthermore, we focused on English data and filtered out non-English tasks. Since our focus is more on evaluating the capabilities of LLMs rather than content moderation, we also removed toxic conversations. To ensure task diversity, we used sentence embeddings from SentenceBERT (Reimers & Gurevych, 2019) to calculate the cosine similarity between queries, discarding those with a high similarity score above 0.9. The threshold is determined by manual inspection. Lastly, to further enhance task diversity, we used a diverse user pool by retaining only the last conversation for each unique device, thus removing tasks from the same user that might require similar underlying skills.
@@
To identify challenging tasks that can distinguish the performance of different LLMs, we used GPT-4-Turbo (OpenAI, 2023), Claude-3-Sonnet, and Opus (Anthropic, 2024) to analyze the required background knowledge and reasoning capabilities for each task. These models assigned a difficulty rating on a five-point scale (from “very easy” to “very hard”). Tasks rated as “very easy” or “easy” by all models were excluded. From the remaining pool, we randomly sampled 1,500 tasks to ensure that the distribution of task categories is similar to the original dataset.
@@
To improve the quality of selected tasks, human annotation was used for quality control. We first used GPT-4-Turbo to summarize the intent of each query. These summaries were then used to help human reviewers remove nonsensical tasks. Finally, we retained 1,024 tasks for WildBench. We also manually reviewed the tasks to ensure that they were challenging and diverse, covering a wide range of task categories. For the checklist questions, we verified that they were clear, interpretable, and relevant to the evaluation of LLM responses.
@@ 2.2 WildBench Statistics
Table 1 compares the statistics of WildBench to existing benchmarks AlpacaEval (Li et al., 2023b; Dubois et al., 2024), MT-Bench (Zheng et al., 2024), and ArenaHard (Li et al., 2024). Among these benchmarks, only ArenaHard and WildBench are sourced from user queries in the wild (“RealUser”), rather than being curated by experts or through crowdsourcing. The difference between ArenaHard and our WildBench is that our data distribution aligns with real users’ task categories, rather than overly focusing on coding and debugging as ArenaHard does.
@@
To enable a fine-grained analysis of LLM capabilities across varied tasks, we categorize the tasks into 12 categories based on previous analysis of ShareGPT queries (Ouyang et al., 2023) and our intent annotation of the tasks. Detailed descriptions about the 12 task categories are shown in Appendix A. The distribution of the task categories is shown in Figure 3. In this figure, we also compare to AlpacaEval and ArenaHard. Notably, WildBench is more balanced compared to AlpacaEval and ArenaHard, which have over 50% of their tasks in Information seeking and Coding & Debugging categories, respectively.
@@ 3 Automatic Evaluation with WildBench
Figure 4: Evaluation framework for WildBench. There are two metrics: WB-Score for individual evaluation and WB-Reward for pairwise evaluation. The checklist is used to guide the evaluation process. The length penalty is used to mitigate the length bias. WB-Reward and WB-Score both have strong correlations with human-based ranking of LLMs on Chatbot Arena.
@@ 3.1 Instance-Specific Checklists
Powerful LLMs have been widely used as judges to evaluate the quality of LLM outputs in many automatic evaluation methods, such as AlpacaEval (Li et al., 2023b). However, even asking humans to judge which of the given two model outputs is better can be subjective and ambiguous. Moreover, such judgements provide limited information about the quality of the models. Without a constant, interpretable, and comprehensive evaluation standard, the results can be noisy and hard to interpret.
To address this issue, we generate a checklist for each test query in WildBench to comprehensively evaluate the responses of different models. The checklist consists of 5-10 questions that are designed to be interpretable and easy to verify. We combine the responses of GPT-4-Turbo and Claude-3-Opus to finalize the checklists, thereby mitigating the bias of using a single LLM as the evaluator. These checklists have been manually reviewed and are used as part of the prompts for LLM judges to evaluate the responses of different models. An example of the checklist can be found in Figure 4. Taking the G20 example in Figure 1, here is a subset of checklist questions for the task:
[…]
@@ 3.2 Pairwise Evaluation with WB-Reward Metric
In Figure 4, we detail the step-by-step evaluation process for pairwise comparison. First, we provide a chain of evaluation questions to guide the LLM judge to analyze the user query and the conversation history. The LLM then evaluates the two responses and also analyze where and why one is better than the other. Finally, we ask the LLM to make a final judgment on which response is better and why. This method is inspired by the evaluation process in human evaluation, where human judges are asked to provide detailed feedback on the quality of the responses before making a final decision. The full evaluation prompt can be found at Appendix D
@@
Using a single baseline model for pairwise evaluation can lead to noisy and biased evaluations. To mitigate this issue, we use three baseline models (GPT-4-Turbo-0429, Claude-3-Haiku, and Llama-2-70B-chat (Touvron et al., 2023)) to compute the rewards for each model. Our metric WB-Reward (Mix) is the average of the rewards from these three baselines on 1024 examples, providing a more robust performance evaluation on WildBench.
@@
Previous studies have shown that LLM judges tend to prefer longer responses (Dubois et al., 2024). To mitigate this bias, we propose a simple and intuitive length penalty method. If the winning response is longer than the losing one by a certain threshold (K characters), we convert Slightly Win/Slightly Lose to a Tie. $K$ can be customized via our leaderboard web-page for personalized configuration. Setting $K=\infty$ will disable the length penalty. We designed this feature to support a more personalized and flexible leaderboard. For example, users who prefer shorter and more concise outputs can set a smaller K if they do not prioritize correlating perfectly with the general human-based model rankings on ChatbotArena. This choice allows for a customized leaderboard experience depending on user preferences.
@@ 3.3 Individual Evaluation with WB-Score Metric
Although pairwise evaluation provides a direct comparison between LLMs, it is usually more expensive and time-consuming than grading each individual LLM generation. To individually evaluate the performance of each model on WildBench, we prompt GPT-4-Turbo to assign a score from 1 to 10 for each model’s response. The full evaluation prompt can be found at Appendix E.
@@
4 Results & Analysis
We analyze the performance of different models on WildBench. We first present the leaderboard analysis, then examine the length bias issue in the evaluation process, and finally discuss the correlation between WildBench-Score and ChatbotArena Elo rating.
Table 2: Evaluation results (subset) of LLMs using WildBench and other benchmarks. Please refer to Figure 6-7 and demo website to view and interact with the full results.